dx = ? x (Vx – 2) A. O ln( Vx – 2) + c B. O2 In(Vx) + c C. O2 In( Vx – 2) + c D. O2 (VX – 2) + c E. O2 In(x - 2) + c
Answer #1

dal I = va (8-2) da a 2a I = √x - 2 = y, dy s zdy 2 dy = 2 in 1 yltc 2 ln | va-2 -21 + toe. 2 bn (Vx-2) to